The AASTEX Macros for Manuscript Preparation

Chris Biemesderfer Revised by Jeannette Barnes, May 1995 American Astronomical Society

1. Introduction

The AAS has developed an author markup package to assist authors in preparing manuscripts that are intended for submission to the AAS-affiliated journals, or to other journals that wish to use it, such as the PASP. The most important aspect of the AASTEX package is that it defines the set of commands (called markup) that can be used to identify the structural elements of papers. When articles are marked up using this set of standard commands, the papers may also be submitted electronically to the editorial offices, with the ultimate goal being electronic generation of the journals themselves.

This guide contains basic instructions for creating manuscripts using the AASTEX markup package, which functions as substyles to the standard L^{A}TEX article style. Authors are expected to be familiar with the editorial requirements of the journals so that they can make appropriate submissions, as well as to have at least a rudimentary knowledge of L^{A}TEX (for instance, knowing how to set up equations using L^{A}TEX commands). A number of useful publications about L^{A}TEX (and TEX) are listed in the reference section of this guide.

Authors who wish to submit papers electronically to the ApJ, the ApJ Letters, or the AJ must use the AASTEX markup package as described in this guide. The package is an option for PASP submissions at this time.

2. The AASTEX Package and Electronic Submissions

Information about acquiring the AASTEX package will be mailed to you electronically if you send an empty email message (some local mailers require a short subject

line) to aastex-instruct@aas.org.

For purposes of producing paper output, the AASTEX package contains a number of L^{A}TEX style files that produce variously formatted pages. There is a “manuscript” style, two “preprint” styles, and two “deluxetable” styles. The manuscript style (aasms4.sty) is used for papers that are submitted to the AAS journals and the PASP for review. The preprint styles render articles in a compact form that may be suitable for distribution among colleagues. The deluxetable styles are used by authors who wish to submit “camera-ready” tables to the journals for publication. See Sections 6 and 7 for further details about the contents of the AASTEX package.

Electronic submission instructions for the various journals are also available electronically. Send an empty email message to the appropriate address below.

                                                            ApJ          apj-instruct@aas.org
                                                            ApJ Letters  apjlett-instruct@aas.org
                                                            AJ           aj-instruct@aas.org
                                                            PASP         pasp-instruct@aas.org

The AAS provides user support for the AASTEX package and for other aspects of electronic submission to the journals. This support is available electronically by send-

                                                      ing e-mail to aastex-help@aas.org.

3. Command Descriptions for Journal Submissions

This section describes the commands in the AASTEX package that an author might enter in a manuscript that is being prepared for electronic submission to one of the journals. The commands will be described in roughly the same order as they would appear in a manuscript. The reader will also find it helpful to examine the sample files (sample1.tex and sample2.tex) that are distributed with the package. Authors are reminded to check the electronic submission procedures and the instructions to authors for the appropriate journals.

3.1. Preamble

In L^{A}TEX manuscripts, the preamble is that portion of the file before the \begin{document} command.

3.1.1. Getting Started

The first piece of markup in the manuscript must declare the overall style of the document. Any commands that appear before this markup will be ignored.

                                                          \documentstyle[12pt,aasms4]{article}

This specifies the main style to be the article style using twelve point fonts, with modifications and additions for the aasms4 substyle. The aasms4 substyle will issue a warning message if the font size is smaller than twelve points, and the size will be set to twelve points; the file will still be processed.

The aasms4 style must be used for the main body of the manuscript. The paper copy produced by this style file will be double-spaced. Any tables included in the main body of the manuscript will also be double-spaced.

Other styles are available, which will be discussed later in this guide (see Section 6).

3.1.2. Defining New Commands

The AASTEX package supports author-defined commands using \newcommand and \def. These definitions should be placed in the document preamble only. The remainder of this section refers to manuscripts submitted electronically to the ApJ or the ApJ Letters; other users of the package may skip to the next section.

Manuscripts submitted electronically to the ApJ or the ApJ Letters are translated to a language called SGML. In general, author-defined commands that are abbreviations are acceptable and can be translated; commands that attempt to define new symbols (e.g., \new-

command{\lte}{\lower 0.5ex\hbox{${}\buildrel<\

over\sim{}$}}) will not translate correctly. In creating new commands authors are encouraged to use \newcommand rather than the \def command.

Authors who need additional symbols not found in L^{A}TEX or AASTEX may use the AMS symbol definitions and fonts that are discussed in Appendix B. If it is not feasible to use the AMS fonts, authors may create their own symbols via \newcommand provided that the symbol is given exactly the same name as in the AMS fonts. This works because author-defined commands with names that match AMS symbol names will be translated as if they were references to the AMS symbols, regardless of the author’s own definition.

Author-defined commands that use any of the commands listed below are unlikely to translate correctly:

\hskip, \vskip, \raise, \raisebox, \lower, \rlap,
\kern, \lineskip, \baselineskip, \char, \mathchar,
\mathcode, \buildref, and \mathrel.

Authors submitting manuscripts electronically to the ApJ and ApJ Letters will have their manuscripts screened by the Production Office at the University of Chicago Press, and a screening report will be returned to the authors along with the referees’ comments. Any definitions or commands that would cause problems in translation to SGML will be identified in the screening report and must be fixed by the author if the electronic file is to be processed successfully.

3.1.3. Editorial Information

A number of markup commands are available for the editorial offices to record the publication history and slug-line data for each manuscript. These commands should be commented out by the author if they appear in a template copy. At this point authors may wish to skip to the next section.

                                                          \received{RECEIPT DATE}
                                                          \revised{REVISION DATE}
                                                          \accepted{ACCEPT DATE}
                                                          \cpright{type}{year}
                                                          \journalid{VOL}{JOURNAL DATE}
                                                          \articleid{START PAGE}{END PAGE}
                                                          \paperid{MANUSCRIPT ID}
                                                          \cpright{TYPE}{YEAR}
                                                          \ccc{CODE}

Copyright information is specified through the commands \cpright and \ccc. The “type” of copyright and the corresponding year are given in \cpright; valid copyright types are as follows.

AAS Copyright has been assigned to

the AAS. ASP Copyright has been assigned to

the ASP. PD The article is in the public

domain. none No copyright is claimed for the

article.

The copyright type is case sensitive, so the type string must be entered exactly as given above. The Copyright Clearing Center code may be given in the \ccc command; the code is taken as regular text, so any special characters, notably ‘$’, must be properly specified.

3.1.4. Short Comment on Title Page

Authors who wish to include a short remark on the title page, such as the name and date of the journal in which an article has been scheduled, may do so with the following command.

                                                          \slugcomment{TEXT}

In the aasms4 style, such comments appear on the title page after the title and authors; in the aaspp4 style, they are placed at the upper right corner of the title page.

3.1.5. Running Heads

Authors are invited to supply running head information. There are generally two different kinds of data in running heads; the left head contains an author list (last names, possibly truncated as “et al.”), while the right head is an abbreviated form of the paper title. This running head information will not be appear on the printed page.

\lefthead{TEXT}
\righthead{TEXT}

Editors and publishers impose varying requirements on the brevity of these data. A good rule of thumb is to limit the list of authors to three, otherwise use et al., and limit the short form of the title to 44 characters. The editors may choose to modify the author-supplied running heads.

3.2. Starting the Main Body

None of the markup that appears in the preamble actually typesets anything; the preamble is only a control section. The author must include a

\begin{document}

command to identify the beginning of the main textual portion of the manuscript.

3.3. Title and Author Information

Title and author identification are by way of the markup commands \title and \author. The author’s principal affiliation is specified with a separate command \affil. Each \author command should be followed by a corresponding \affil.

\title{LUCID TEXT}
\author{NAME(S)}
\affil{AFFILIATION}
\authoraddr{ADDRESS}
\and

Line breaks are permitted in the title if the author wishes to specify them with the \\ command. Long titles will be broken automatically, so the \\ markup is not required. If the title is explicitly broken over several lines, the preferred style for titles in AAS and ASP journals is the so-called “inverted pyramid” style. In this style, the longest line is the first (or top) line, and each succeeding line is shorter. The text of the title should be entered in mixed case; it will be converted to uppercase by the publisher. Footnotes are permissible in titles; be careful to ensure that alternate affiliations (see below) are properly numbered if a footnote to the title is specified.

Authors’ names are given in \author commands, and should be entered in mixed case. Names that appear together in the author list for authors who have the same primary affiliation should be specified in a single \author command. Each author group (\author command) should be followed by an \affil command, giving the principal affiliation of that author. Physical and postal address information for the institution specified may be included with its name. The address can be broken over several lines by using the \\ command to indicate the line breaks. Usually, however, postal information will fit on one line. When there is more than one \author command, the last one should be preceded by an \and command.

When there is a lengthy author list, all authors’ names may be specified in a single \author command, with affiliations specified using the \altaffilmark mechanism described below. In this case, no \affil commands would be used, and the affiliations would all be listed in a footnote block at the bottom of the title page. The style file performs this formatting.

Postal addresses for individual authors may be specified in \authoraddr commands. This command does not produce any formatted text in most AASTEX styles and can be used to specify the corresponding address of the first author for purposes of editorial communication.

Authors often have affiliations in addition to their principal employer, and these are specified with the

\altaffilmark and \altaffiltext commands. These
behave like the \footnotemark and \footnotetext com-

mands of L^{A}TEX, except there are no optional arguments in the AASTEX commands. \altaffilmark is appended to author’s names in the \author lists and generates superscript identification numbers. The text for the individual alternate affiliations is generated by

the \altaffiltext command.
     \altaffilmark{KEY NUMBER(S)}
     \altaffiltext{NUMERICAL KEY}{TEXT}

It is up to the author to make sure that \altaffilmark numbers attached to authors’ names correspond to the correct alternate affiliation, i.e., that each KEY NUM-BER matches the NUMERICAL KEY for the corresponding TEXT.

3.4. Abstract

The paper abstract should be enclosed in an abstract environment.

     \begin{abstract}

abstract text

     \end{abstract}

3.5. Keywords

Keywords, subject headings, etc., are accommodated as a single piece of text.

\keywords{TEXT}

If authors supply keywords they must be delimited by the appropriate punctuation required by the journal. Keywords should be specified in alphabetical order. The \keywords command supplies the proper leading text (“Keywords:”, “Subject headings:”, etc.), according to the style.

The list of keywords used by the ApJ and the ApJ Letters are in the file keywords.apj in the directory

pubs/aastex-misc on the anonymous ftp node aas.org.

3.6. Comments to Editors

Authors may make notes or comments to the editors with the \notetoeditor{TEXT} command. This command will behave like a footnote and appear on the bottom of the page, e.g., ^{E1}: NOTE TO THE EDITOR: <TEXT>. Output to the printed page is produced with the

aasms4.sty file only.
\notetoeditor{TEXT}

3.7. Sections

The article style for AASTEX manuscripts supports four levels of sectioning.

\section{HEADING}
\subsection{HEADING}
\subsubsection{HEADING}
\paragraph{HEADING}

Section headings should be given in mixed case. Note that these commands delimit sections by marking the beginning of each section; there are not separate commands to identify the ends.

3.8. Figure and Table Placement

Generally figures and tables are not “placed” in the text of the document where an author would like them to appear physically, but rather follow the main body of the text. However, authors may indicate to the editors the preferred placement of these items in the text by use of the \place* commands.

\placetable{KEY}
\placefigure{KEY}

The \place*{KEY} commands are similar to the \ref command in LaTEX and require corresponding \label commands to link them to the proper elements for assisting the editors. The \label{KEY} command may be included in a \figcaption command used to produce figure captions on the figure caption page or in a \caption command used in the figure environment (see Sections 3.15 and 5); in a \tablecaption command in the deluxetable environment (see Section 3.13.1), or in a \caption command in a LaTEX table environment.

On the printed page the \place* commands will print a short message to the editor about figure and table placements when used with the aasms4.sty file. Nothing will be printed with the other style files.

3.9. Acknowledgments

In addition, AASTEX manuscript styles support an

\acknowledgments section.
     \acknowledgments

In the AASTEX substyles, acknowledgments are set off from the concluding main body text simply by vertical space (no heading or type size change).

3.10. Appendices

When one or more appendices are needed in a paper, the end of the main body text must be marked.

     \appendix

Note that the \appendix command has no arguments; sections in the appendix must be headed with \section commands containing the section headings, as described earlier. The \appendix command takes care of a number of internal housekeeping concerns, such as identifying sections with letters instead of numerals, resetting the equation counter, etc.

3.11. Citations

Two options are available for marking citations within a manuscript. The \cite command is used in conjunc-

tion with thebibliography.
     \cite{KEY}

L^{A}TEX’s thebibliography environment allows authors to identify references symbolically using unique keys (the author makes these up). If this mechanism is used, the KEY given in the \cite command must correspond to a KEY given in a \bibitem command in the thebibliography environment. The LABEL in the \bibitem command is used as the citation in the text. See Section 3.14.1.

The conventional method used by authors to manage the citations and reference list in a paper has often been a manual one. These authors can use the references environment and \markcite command.

\markcite{KEY}

The KEY supplied in the \markcite command must match the key used by the \reference command for the corresponding reference in the references environment. The \markcite command is only used to point to the proper reference in the reference list—the actual citation must be supplied by the author as part of the running text. See Section 3.14.2.

The use of the references environment and the \markcite command has changed with the v4.0 release of the AASTEX package and is not backward or forward compatible with v3.0.

3.12. Equations

Displayed equations can be typeset in many ways using the standard displayed math environments of L^{A}TEX; these three are probably of greatest use.

\begin{displaymath}
\end{displaymath}
\begin{equation}
\end{equation}
\begin{eqnarray}
\end{eqnarray}

The displaymath environment will break out a single, unnumbered formula. The same formula will appear the same if it is set in an equation environment, except it will be autonumbered by L^{A}TEX. In order to set several formulæ in which vertical alignment is required or to display a long equation across multiple lines, the eqnarray environment should be used.

Authors occasionally wish to group related equations together and identify them with letters appended to the same equation number, as opposed to having each with a separate numeral. When this is desired, such related equations should still be set in equation or eqnarray environments (whichever is appropriate may be used), and this grouping should be placed within a mathletters environment.

\begin{mathletters}
equation or eqnarray environment(s)
\end{mathletters}

It is possible to override L^{A}TEX’s automatic numbering within equation or eqnarray environments.

     \eqnum{TEXT}

When \eqnum is specified inside an equation environment, or on a particular equation within an eqnarray, the text supplied as an argument to \eqnum is used as the equation identifier. L^{A}TEX’s equation counter is not incremented when \eqnum is used. \eqnum must be used inside the environment.

When unnumbered equations are desired, authors can use either the displaymath environment (for single displayed equations) or place a \nonumber command before the equation delimiter (\\) in a particular equation in an eqnarray. L^{A}TEX’s equation counter is not incremented when \nonumber is used.

If, as a consequence of the use of \eqnum or \nonumber, L^{A}TEX’s equation counter gets out of the author’s intended sequence, the counter may be reset to a particular value.

     \setcounter{equation}{NUMBER}

The equation counter should be set to the number corresponding to the last equation that was formatted, so it is most appropriate for this command to appear immediately after an equation or eqnarray environment ends. \setcounter{equation} must be used outside the math environments.

3.13. Tables

There is support in the AASTEX package for tables via two mechanisms: L^{A}TEX’s standard table and tabular environments, and a deluxetable environment that facilitates the formatting of tabular (particularly lengthy tabular) material. Tables may be marked up using either mechanism although the deluxetable environment is preferred. Authors should not use the L^{A}TEX tabbing environment or the \hline command for electronic submissions.

L^{A}TEX permits the preparation of fairly complex tables with arbitrary spacing, straddle heads and rules, and the like. Authors who need to specify complicated column headings and so forth are advised to consult the L^{A}TEX manual (Lamport 1985) for details. Most of the capabilities are applicable to AASTEX’s deluxetable environment as well as L^{A}TEX’s tabular.

Authors should be aware that all tables produced using the aasms4 style will be double-spaced. If singlespaced, or “camera-ready” tables are required, then an alternate style file must be used for the tables only (see Section 4).

3.13.1. The deluxetable Environment

This section describes the use of the deluxetable environment. There are several desiderata that are somewhat above and beyond LaTeX’s tabular environment that facilitate the formatting of such tables. Among these are breaking long tables across pages, using footnotes in a table, specifying comments and references for tables, etc.

The deluxetable environment is delimited by L^{A}TEX’s familiar \begin and \end constructs.

\begin{deluxetable}{COLS}
\end{deluxetable}

COLS specifies the justification for each column. One of the letters ‘l’, ‘c’, or ‘r’ is given for each column, indicating left, center, or right justification. Authors are referred to the L^{A}TEX manual (Lamport 1985) for further information.

There are several items in a deluxetable environment that must be given before the data for the table.

\small
\footnotesize
\scriptsize
\tablewidth{DIMEN}
\tablenum{TEXT}
\tablecaption{TEXT \label{KEY}}

If a table is too wide for the printed page, it is permissible to change the font size within the deluxetable environment using the \small (11pt), the \footnotesize (10pt) or the \scriptsize (8pt) command.

The width of a deluxetable is defined by \tablewidth; the default width, if this command is omitted, is the width of the body text. The table can be set to its natural width by specifying a DIMEN of 0pt. Long tables may have a natural width that is different for each page. The natural width for each page will be printed to the log file during processing; authors may then define a fixed table width based on this information giving the tables a more uniform appearance across the pages.

It is possible to override L^{A}TEX’s automatic numbering within the deluxetable environment. When \tablenum is specified inside a deluxetable environment, the text supplied as an argument to \tablenum is used as the table identifier. L^{A}TEX’s equation counter is not incremented when \tablenum is used. \tablenum must be used inside the deluxetable environment and before the

\tablecaption command.

The caption (actually, the title) of the table is specified in \tablecaption. The intent is for the text of \tablecaption to be brief; explanatory notes may be specified in the end notes to the table (\tablecomments, see below). If the caption does not appear centered on the table after processing, then specify the width of the table explicitly in the \tablewidth command and rerun L^{A}TEX on the file.

The KEY in the label command should be the same KEY used in the \placetable{KEY} command used to call out the preferred location of the table in the text. The label command, if used, must be included in the

\tablecaption command.
     \tablehead{TEXT}
     \colhead{HEADING}

Column headings are specified within \tablehead. Within \tablehead, each column heading can be given in a \colhead, which will ensure that the heading is centered on the natural width of the column; this is the typical disposition of column headings, and the use of \colhead is encouraged. There should be a heading for each column, so there should be as many \colhead commands in the \tablehead as there are data columns. If more complicated column headings are required, any valid tabular commands that constitute a proper head line for the table may be used. Consult the L^{A}TEX manual (Lamport 1985) for details about using the tabular environment to prepare tables.

The \tablecolumns{NUM} command is necessary if the author has multiline column headings produced by \tablehead or other L^{A}TEXcommands and is using either the \cutinhead or \sidehead markup (see below). The NUM is set to the true number of columns in the table, and the command must come before the \startdata command.

     \tablecolumns{NUM}

See the complext.tex file in the AASTEX package for an example of a more complex table format using the

\tablecolumn command.

Column alignment within the data columns can be adjusted with the TEX \phantom{STRING} command, where STRING can be any character representation, e.g., \phantom{$\arcmin$}. A blank character of width STRING is then inserted in the table. Three characters have been predefined for this purpose as well as a shorter version of the generic phantom command:

\phn  phantom numeral 0-9
\phd  phantom decimal point
\phs  phantom plus or minus sign
\phm{STRING} generic phantom command

It is possible that a complicated table heading will overflow the vertical space allotted for the table heading. The fraction of the page allocated for the table heading may be changed with \tableheadfrac. The NUM argument to \tableheadfrac should be the decimal fraction of the page used for heading information. The default value is 0.1, meaning that 10% of the page height is reserved for the table heading. It should rarely be necessary to change this value.

     \tableheadfrac{NUM}

After the table title and column headings are specified, data lines can be entered. The data lines are within two “data” commands. Data elements within a row of the table are separated with & (ampersand) characters. The end of each row is indicated with \nl.

     \startdata

data lines

     \enddata
     \nl
     \tablevspace{DIMEN}

Extra vertical space can be inserted between rows with a \tablevspace command; the argument is a dimension and may be specified in any units that are legitimate in L^{A}TEX.

     \tablebreak
     \nodata

If a page break needs to be forced in a deluxetable, \tablebreak should be used instead of \nl. This is sometimes necessary when several rows of data are associated with a single object or item; such logical groupings should not be broken across pages, and \tablebreak can be used to ensure that the page breaks are reasonable in these cases.

The journals often require that elements for which there are no data be explicitly marked. This is to differentiate such elements from blank elements, which are frequently interpreted as implicitly repeating the entry in the corresponding element in the row preceding. Data elements for which there are no data should contain a \nodata command; an appropriate symbol will be placed in that data element.

Within the deluxetable body, two kinds of “specialty” heads are recognized. A cut-in head is a piece of text that is centered on the table width; it is spaced above and below from the data rows that precede and follow it, and there may be rules associated, depending on the journal or manuscript style. All of these formatting particulars are managed by the style files. The author need only specify the text to be centered with a \cutinhead command. Similarly, a side head is a piece of text that is left-justified.

                                                          \cutinhead{TEXT}
                                                          \sidehead{TEXT}

Table footnotes (more properly, table endnotes) may be used in the deluxetable environment; their use is described in detail in the section on Table Footnotes, below. These notes must follow the \enddata command. The markup commands for such endnotes are as follows.

                                                          \enddata
                                                          \tablenotemark{KEY LETTER(S)}
                                                          \tablenotetext{ALPHA KEY}{TEXT}
                                                          \tablecomment{TEXT}
                                                          \tablerefs{TEXT}

3.13.2. The table Environment

Tables may be marked and composed using the standard L^{A}TEX tools for tables, although the use of the deluxetable environment is encouraged. Tables should appear in table environments.

                                                          \begin{table}
                                                          \end{table}

The table environment encloses not only the tabular material but also any title (caption) or footnote information associated with the table.

Titles or captions for tables are indicated with

                                                          \caption{TEXT  \label{KEY}}

Tables will be identified with arabic numerals, e.g., “Table 2”; the identifying text, including the number, is generated automatically by \caption within the table environment. The KEY in the label command should correspond to the KEY used in the \placetable{KEY} command used to identify the placement of the table in the text.

Tabular information is typeset within the tabular environment:

                                                          \begin{tabular}{COLS}
                                                          \end{tabular}

where COLS specifies the justification for each column. One of the letters ‘l’, ‘c’, or ‘r’ is given for each column, indicating left, center, or right justification. Consult the L^{A}TEX manual (Lamport 1985) for details about using the tabular environment to prepare tables. A tabular table must appear within a table environment. And there should be only one tabular table per table environment. If the journal requests manuscripts with only one table per page, the author may need to insert a \clearpage command after especially short tables.

There is a \tableline command for use in tabular environments.

     \tableline

This command produces the horizontal rule(s) between the column headings and the body of the table. Authors are discouraged from using any \hlines themselves; vertical rules are typically forbidden by editorial preference.

It is possible to override L^{A}TEX’s automatic numbering within the table environment.

     \tablenum{TEXT}

When \tablenum is specified inside a table environment, the text supplied as an argument to \tablenum is used as the table identifier. L^{A}TEX’s equation counter is not incremented when \tablenum is used. \tablenum must be used inside the table environment.

3.13.3. Table Footnotes

AASTEX supports footnotes (endnotes) that are associated with tables; this support applies to both the deluxetable environment and the standard L^{A}TEX table environment. Footnotes for tables are usually identified by lowercase letters rather than numbers. Marking and assigning associated text is achieved with the

\tablenotemark and \tablenotetext commands, in

which the note identifier is required (cf. \altaffilmark

and \altaffiltext). The \tablenotetext must be

specified before the enclosing \end{table}, since the text of notes to tables are displayed by that command.

     \tablenotemark{KEY LETTER(S)}
     \tablenotetext{ALPHA KEY}{TEXT}

Note that the KEY LETTER should be the same as the ALPHA KEY for the corresponding TEXT. It is the responsibility of the author to make the correspondence correct.

AASTEX also supports special kinds of table endnotes. Sometimes authors tabulate materials which have corresponding references, and it may be desirable to associate these references with the table rather than (or in addition to) the formal reference list. Occasionally, authors wish to append a short paragraph of explanatory notes that pertain to the entire table, but which are different than the caption.

                                                          \tablerefs{REFERENCE LIST}
                                                          \tablecomments{TEXT}

The table endnotes are coupled to the table in which they occur, rather than being associated with a particular page, and they are printed with the table (relatively close to the caption) instead of appearing at the extreme bottom of the page. This is done to ensure that the notes wind up on the same page as the table, since tables are floats and can migrate from one page to another.

3.14. Bibliography

Observe that the bibliographic data supplied by the author must conform to the standards of the journal. We have elected not to burden authors with tedious markup commands to delimit the bibliographic fields because many of the journals we have targeted in this project have agreed to reduce typographic overhead (bolding, italicizing, etc.) in reference lists (Abt 1990). It is the responsibility of the author to arrange these fields in the proper order with the correct punctuation; the information will be typeset as is, i.e., in roman with no size or style changes.

An identifier denoting that the article was prepared with the AASTEX package appears as a footnote on the last page of references.

Authors has two options for marking citations in the text and providing the associated reference lists.

3.14.1. The thebibliography Environment

Authors are encouraged to use the semantics of L^{A}TEX’s thebibliography environment, marking citations with \cite and associating references with them via \bibitem. The \cite-\bibitem mechanism associates citations and references symbolically while maintaining proper citation syntax within the paper. The author must create the citation label for each reference in proper journal format in the \bibitem command.

                                                          \begin{thebibliography}{DUM}
                                                          \bibitem[LABEL]{KEY} bibliographic data
                                                             .
                                                             .
                                                          \end{thebibliography}

where LABEL must adhere to journal standards, e.g., “Abt 1986”. The value of LABEL is then inserted into the body of the text at the place of the \cite command during the processing. The argument DUM to thebibliography environment is not used by the AASTEX package. Note that it is not possible to use \bibitems within the references environment, nor will \cite commands work properly in the main body if \bibitems are not properly specified. This technique can be a bit tricky, and there are limitations on the way that the citation LABEL is formatted.

Citation management can be complex, and systems have been developed to assist authors in preparing bibliographies. The program that manages references within the TEX family is called BIBTEX, and it is designed to work in conjunction with the citation and reference list capabilities of L^{A}TEX. Authors are advised to consult the L^{A}TEX manual (Lamport 1985). The AASTEX package has no additional support for BIBTEX.

3.14.2. The references Environment

Many authors are comfortable with the standard process of entering citations in the proper format directly in the body of an article and then organizing the reference list themselves. Such authors would use \markcite{KEY} in the text before the citation, e.g., ...as discussed by

Marlow \markcite{a3} (1993), and then employ the

references environment for the reference list, where each reference is preceded by a \reference{KEY} com-

mand, e.g., \reference{a3} Marlow, R. 1993, \nat,

445, 12. The \markcite{KEY} command indicates in the text where a certain reference in the reference list is used; the author must still supply the correct citation. The references environment simply sets off the list of references and adjusts spacing parameters.

\begin{references}

\reference{KEY} bibliographic data

.
.
\end{references}

The implementation of the \markcite-\reference markup has changed with the v4.0 release of the AASTEX package. Its use is not backward or forward compatible with v3.0.

3.14.3. Abbreviations for Journals

There are markup commands for many of the oftreferenced journals so that authors may use the markup names as a shorthand rather than having to look up a particular journal’s specific abbreviation. In principle, all the journals should be using the same abbreviations as well, but it is fair to anticipate some changes in the specific abbreviations before a system is finally settled on. As long as these commands are kept up to date, authors need not be concerned about such editorial preferences and changes.

\araa  Annual Review of Astronomy

and Astrophysics

\ao  Applied Optics
\aj  Astronomical Journal
\azh  Astronomicheskii Zhurnal
\aap  Astronomy and Astrophysics
\aapr  Astronomy and Astrophysics Reviews
\apj  Astrophysical Journal
\apjl  , Letters to the Editor
\apjs  , Supplement Series
\aplett  Astrophysics Letters
\apspr  Astrophysics Space Physics Research
\apss  Astrophysics and Space Science
\aaps  , Supplement Series
\baas  Bulletin of the AAS
\bain  Bulletin Astronomical Inst. Netherlands
\gca  Geochimica Cosmochimica Acta
\grl  Geophysics Research Letters
\iaucirc IAU Circular
\jcp  Journal of Chemical Physics
\jgr  Journal of Geophysics Research
\jrasc  Journal of the RAS of Canada
\memras  Memoirs of the RAS
\mnras  Monthly Notices of the RAS
\nat  Nature
\nphysa  Nuclear Physics A
\physscr Physica Scripta
\pra  Physical Review A: General Physics
\prb  Physical Review B: Solid State
\prc  Physical Review C:
\prd  Physical Review D:
\pre  Physical Review E:
\prl  Physical Review Letters
\physrep Physics Reports
\planss  Planetary Space Science
\procspie Proceedings of the SPIE
\pasj  Publications of the ASJ
\pasp  Publications of the ASP
\qjras  Quarterly Journal of the RAS
\skytel  Sky and Telescope
\solphys Solar Physics

3.15. Figure Captions

Figure captions or legends are included in the manuscript on a separate page using the \figcaption command for each figure in the paper.

    \newpage
    \figcaption[FILENAME]{TEXT \label{KEY}}

The optional argument, FILENAME, can be used to identify the PostScript file for the corresponding figure, e.g., graph.eps; TEXT refers to the caption for that figure.

The KEY in the label command refers to the KEY in the \place*{KEY} command used to identify the placement of the figure in the manuscript (see Section 3.8).

When the \figcaption command is used, the figure identification, e.g., “Figure 1,” is generated automatically by the command itself.

Some journals may accept figures placed in the figure environment (see Section 5) if the manuscript is being submitted electronically. Please check the appropriate electronic submission instructions for your target journal.

3.16. Miscellaneous

When discussing atomic species, ionization levels can be indicated with the following command.

     \ion{ELEMENT}{LEVEL}

The ionization state is specified as the second argument, and should be given as a numeral. For example, Ca III is specified by typing \ion{Ca}{3}.

AASTEX contains two commands that permit authors to specify alternate forms for fractions. Authors submitting manuscripts electronically will generally find it unnecessary to use any markup other than the standard L^{A}TEX \frac.

     \case{NUM}{DENOM}
     \slantfrac{NUM}{DENOM}

L^{A}TEX will set fractions in displayed math as built-up fractions; it is sometimes desirable to use case fractions in displayed equations. In such instances, one should use \case rather than \frac. Occasionally, authors wish to typeset fractions with a solidus but in which the type size is reduced and the numerals are oriented diagonally. In this case, \slantfrac should be used instead of \frac. Note that this is different from a shilled fraction, which authors can produce without any special markup.

Built-up \frac\{1\}\{2\}  \frac{1}{2}
Case  \case\{1\}\{2\}  \frac{1}{2}
Slant  \slantfrac{1}{2} ^{1}/_{2}
Shilled 1/2  1/2

The AASTEX package also contains a collection of assorted macros for special symbols (or abbreviations) that authors tend to work out for themselves anyway. Some of the definitions come from the A&A package (Springer 1989); some are contributions from individuals. We have tried to select a tractable number that were useful and also somewhat difficult to get right because fussy kerning or some such is required.

                                                        \arcdeg         ^{∘}  \sq  ⊓ ⊔
                                                        \sun  ⊙  \earth  ⊕
                                                        \arcmin         ′  \arcsec    ′′
                                                        \fd  .^{d}  \fh  .^{h}
                                                        \fm  .^{m}  \fs  .^{s}
                                                        \fdg  .^{∘}  \farcm  .′
                                                        \farcs  .′′  \fp  .^{p}
                                                        \micron  μm
                                                        \onehalf         ^{1}/_{2}  \ubvr  UBVR
                                                        \onethird        ^{1}/_{3}  \ub  U −B
                                                        \twothirds       ^{2}/_{3}  \bv  B−V
                                                        \onequarter      ^{1}/_{4}  \vr  V −R
                                                        \threequarters   ^{3}/_{4}  \ur  U −R
                                                        $\lesssim$  ∼ <  $\gtrsim$ ∼ >

Most of these commands can be used in running text as well as when setting mathematical expressions. \lesssim and \gtrsim can only be used in math mode, which is sensible since they are relations. It is possible to use

                                                      \earth and \sun as subscripts, e.g., $1.4 M_{\sun}$

yields 1.4M⊙.

3.17. Concluding the File

The last command in the electronic manuscript file should be the

                                                          \end{document}

command, which appears after all the material in the paper. This command directs the formatter to perform assorted termination activities and finish processing.

4. Camera-ready Tables

Long tables that are prepared as “camera-ready” and submitted on paper will require the use of a deluxetable substyle file.

There are special L^{A}TEX substyle options within the AASTEX package which authors can use to produce deluxetable output suitable for photographic reproduction in specific AAS and ASP journals; these are the aj_pt4 and apjpt4 substyles.

                                                          \documentstyle[aj_pt4]{article}
                                                          \documentstyle[apjpt4]{article}

When tabular material is prepared with one of these style files, the pages produced will be suitable for submission to the journal as camera-ready “art.” These special styles are intended for producing deluxetables only; manuscripts should not be formatted with the aj_pt4 or the apjpt4 substyle. See the example files

samp2tbl.tex, samptbls.tex, and complext.tex in the

AASTEX package.

In general, the same markup commands apply to these tables as were described earlier for the deluxetable environment (see Section 3.13.1). The commands \small,

\footnotesize, and \scriptsize are not permitted,

however.

If the table is too wide for the page, the \ptlandscape command may be used to format the table in landscape mode; this command must be issued before the \begin{deluxetable} command. The author must remember to print the table using the landscape option in the dvi print program, e.g., dvips -t landscape.

These special substyle options for tables, aj pt4.sty and apjpt4.sty, may also be used to produce tables that are not meant to be “camera-ready” but are too wide to fit on a page even after using the techniques described in Section 3.13.1. The author could use the \ptlandscape option but double-space the table as required for normal manuscript submission.

     \ptlandscape
     \begin{deluxetable}
     \doublespace

Authors may reference these external tables from within their main manuscript file with the \ref{KEY} command by using the \dummytable command within the table environment as part of their main manuscript file.

     \begin{table}
     \dummytable\label{KEY}
     \end{table}

5. Figures

At this time, the most widely used means of including figures or other types of nontextual data in electronic manuscripts is to generate such files as PostScript.^{1}

For manuscripts that are submitted electronically the figures or PostScript files are generally sent as individual files to the editors and are not included as part of the main body of text. The legends or captions for these files are included in the main body of the text, however, on a separate page with specific formats (see Section 3.15). Authors are referred to the submission instructions for the appropriate journals if PostScript figures are to be submitted along with an electronically submitted manuscript.

If the author wishes the PostScript figures to be included in the pages being produced, it is necessary to use the dvips program. This program is available via anonymous ftp from labrea.stanford.edu. It is also necessary that the graphics files being included conform to the Encapsulated PostScript standard (Adobe 1990). Encapsulated PostScript will be referred to as ‘EPS’ in what follows.

There are several commands associated with including EPS files in AASTEX manuscripts, and these should be placed within a figure environment.

                                                          \begin{figure}
                                                          \figurenum{TEXT}
                                                          \epsscale{NUM}
                                                          \plotone{EPSFILE}
                                                          \plottwo{EPSFILE}{EPSFILE}
                                                          \plotfiddle{EPSFILE}{VSIZE}
                                                              {ROT}{HSF}{VSF}{HTRANS}{VTRANS}
                                                          \caption{TEXT}
                                                          \end{figure}

When \figurenum is specified inside a figure environment, the text supplied as an argument to \figurenum is used as the figure identifier. L^{A}TEX’s figure counter is not incremented when \figurenum is used. \figurenum must be used inside the figure environment.

\plotone inserts the graphic in the named EPSFILE, scaled (in both dimensions) so that the horizontal dimension fits in the body text width; the vertical dimension is scaled to maintain the aspect ratio. \plottwo inserts two plots next to each other. Scale factors are determined automatically from information in the EPS file.

The automatic scaling may be overridden with the command \epsscale{NUM}, where NUM is in decimal units, i.e., 0.80.

For the adventuresome, there is also \plotfiddle, which can be used to override the automatic figure placement, scaling, and rotation. When this method is used, the author must specify the vertical space allotment for the graphic (see VSIZE below). The scaling and placement of the figure are controlled by these parameters:

1PostScript is a registered trademark of Adobe Systems Incorporated.

VSIZE vertical white space to allow for plot (L^{A}TEX dimension)

ROT  rotation angle (degrees)
HSF  horizontal scale (percentage)
VSF  vertical scale (percentage)
HTRANS horizontal translation (PS points)
VTRANS vertical translation (PS points)

PostScript points are 1/72 inches, so an HTRANS of 72 moves the graphic 1 inch to the right. The PostScript reference point is the lower left corner of the page, so a VTRANS of 72 moves the graphic up 1 inch.

There is an upper limit of seven \begin...\end figure environments per page.

Footnotes are not supported for figures.

6. Alternate Style Options

It is not the objective of the AASTEX project to develop L^{A}TEX styles that produce pages that mimic the appearance of specific journals. Since this is the case, we have chosen to provide several format options within the preprint styles themselves so that output format can be varied. The aaspp4 substyle offers a single-spaced alternative for manuscripts, for instance. The aaspp4 substyle is similar to the manuscript style, but it is singlespaced, and it is possible to use a smaller typeface. For institutional purposes, it may be preferable for preprints to be set in two columns, have running heads, etc. The aas2pp4 substyle may be used instead to produce twocolumn pages. This style can be used as is, but it can also serve as a point of departure for L^{A}TEX style writers at institutions that want preprints of this general nature. These two style options are discussed further below.

The preparation of “plano tables” (camera-ready, or planographic) is possible by using one of the style files aj_pt4 or apjpt4. These styles produce tables in formats suitable for the AJ and the ApJ, respectively; the aj_pt4 style could be used to create camera-ready tables for the PASP as well. These styles are intended primarily for the production of tables that the author wishes to submit camera-ready; they are not to be used for producing manuscript pages.

The eqsecnum style file can be used to modify the way equations are numbered. Normally, equations are just numbered sequentially through the entire paper. If eqsecnum appears in the documentstyle command, equation numbers will be sequential through each section and will be formatted “(sec-eqn),” where “sec” is the current section number and “eqn” is the number of the equation within that section. The eqsecnum option can be used with the manuscript or preprint substyle files. For example,

                                                          \documentstyle[11pt,eqsecnum,aaspp4]
                                                              {article}

A flushrt.sty file option is available for authors that prefer to have their margins left and right justified. The aas2pp4 style option is flush right by default, but the aasms4 and aaspp4 style options are ragged right by default.

                                                          \documentstyle[11pt,eqsecnum,aaspp4,flushrt]
                                                              {article}

6.1. Preprint format

A single-column preprint format can be specified with the aaspp4 substyle option.

                                                          \documentstyle[11pt,aaspp4]{article}

The size of the typeface used is under author control by way of L^{A}TEX’s NNpt article substyles (where NN is 10, 11, or 12). Use of 10 point type is not recommended with the aaspp4 style.

Authors may wish to adjust vertical spacing within a preprint, for instance, double-spacing text while singlespacing tables. Authors who want to alternate between single and double spacing in the manuscript may use the following commands.

                                                          \singlespace
                                                          \doublespace

\singlespace sets the vertical spacing to a smaller value, while \doublespace causes double-spacing.

6.2. Two-column format

The aas2pp4 substyle has the principal function of setting up two-column output.

                                                          \documentstyle[aas2pp4]{article}

Although it is quite obvious, it is important to remember that text lines are considerably shorter when two of them are typeset side by side on a page. Long equations, wide tables and figures, and the like, may not typeset in this format without some adjustments. The expenditure of great effort to adapt copy and markup for twocolumn pages is probably counterproductive. Remember that the main goal of this package at this point is to produce “correct” draft (or referee) format pages; it is the responsibility of the editors and publishers to produce publication format papers for the journals.

The aas2pp4 substyle does not impose a format for the article’s front matter, although there is often merit in setting the title, author, abstract, and keyword material on a separate page at full text width. The author may supply the \twocolumn command wherever desired.

     \twocolumn

Note that the two-column format begins at the point \twocolumn appears in the text, and if that point is before the front matter, title and author and so forth will be typeset in two-column mode along with the rest of the paper; that is how this manual is prepared, for example. For purposes of producing “pretty” output, it is probably desirable to put the \twocolumn command after the abstract and keywords, just before the body of the paper. If \twocolumn is not specified explicitly in aas2pp4 style documents, the introductory material of the paper will be set in one-column mode; the first \section command (presumably demarcating the beginning of the main body of the article) will engage the two-column mode.

7. Additional documentation

The preceding detailed explanation of the markup commands in this package has certain merit, but many authors will prefer to examine the sample papers that are included with the style files. The files of interest are described below.

The file sample1.tex is a paper prepared with the AASTEX package utilizing a minimal amount of markup. The most salient thing to observe about this example is that, apart from the document style declarations, no formatting instructions are given in the file.

A more comprehensive example requiring nearly all of the capabilities of the package (in terms of markup as well as formatting) is in sample2.tex. This file is annotated with comments that describe the purpose of most of the markup. sample2.tex has three tables: one marked up using the deluxetable environment; another table using the L^{A}TEX table environment; and the third table, samp2tbl.tex, produced as a camera-ready table using either the apjpt4 or aj_pt4 style files.

A more complex example of the deluxetable environment is presented as a camera-ready table in the file

complext.tex.

A set of three tables prepared as “deluxe” tables are contained in the files table1.tex, table2.tex, and table3.tex. They can be formatted by running L^{A}TEX on the samptbls.tex file, into which they are included. The document style in samptbls.tex is apjpt4 in the distribution, but this can be changed to aj_pt4 to see the effects of differing requirements among journals. Page breaks are explicitly indicated in table1.tex, and they are set for the ApJ style; due to the different type size used in the AJ, the pages break somewhat erratically when the aj_pt4 style is used. It is left as an exercise to the reader to make sense of this.

This user guide (aastex.tex) is also marked up with the AASTEX package, although it is not exemplary as a scientific paper.

A number of the markup commands described in the preceding sections are standard L^{A}TEX commands, and the reader who is unfamiliar with their syntax is referred to the L^{A}TEX manual (Lamport 1985) or another guide by Hahn (Hahn 1993) for details. A crib sheet listing all the L^{A}TEX commands (and some pertinent plain TEX commands) with short descriptions of each is published by the TEX Users Group (Botway and Biemesderfer 1989). Authors who wish to know the ins and outs of TEX itself should read the TEXbook (Knuth 1984), probably more than once. There is a good deal of information about typography in general in this source. Many details of mathematical typography are discussed in a book by Swanson (1971).

REFERENCES

Abt, H. 1990, ApJ, 357, 1 (editorial)

Adobe Systems, Inc. 1990, PostScript Language Reference Manual, Appendix H (Reading, MA: Addison-Wesley)

Biemesderfer, C., and Hanisch R. 1989, BAAS, 21, 780 Botway, L., and Biemesderfer, C. 1989, L^{A}TEX Command Summary (Providence, RI: TEX Users Group) Hahn, J. 1993, L^{A}TEX for Everyone (Englewood Cliffs,

NJ: Prentice-Hall) Hanisch, R., and Biemesderfer, C. 1990, BAAS, 22, 829 Knuth, D. 1984, The TEXbook (Reading, MA: Addison-

Wesley) Lamport, L. 1985, L^{A}TEX: A Document Preparation Sys-

tem (Reading, MA: Addison-Wesley) Springer-Verlag. 1989, Springer-Verlag TEX AA macro

package 1989 (Springer: Heidelberg) Springer-Verlag. 1990, Springer-Verlag L^{A}TEX AA macro

package 1990 (Springer: Heidelberg) Swanson, E. 1979, Mathematics into Type (Providence,

RI: American Mathematical Society)

This 2-column preprint was prepared with the AAS L^{A}TEX macros v4.0.

A. Some Helpful TEX Hints

In the sections that follow, we review some essential procedures that must be followed when preparing TEX input.

A.1. Running text

Printing is different from typewriting, and TEX is different from other word processing tools. This section consists of reminders (admonitions) about things that require special attention so that TEX can format the input properly.

• The ends of words and sentences are marked by white space, and it does not matter how many spaces are typed; one is as good as 100. TEX treats the end of a line in the input file as a space.

• Paragraphs are separated by blank lines. Do not hyphenate words in the input file; TEX takes care of hyphenation automatically. Continue to hyphenate modifiers within a line of text, e.g., “authorprepared copy.”

• Quotation marks should be typed as pairs of opening and closing single quotes, e.g., ``quoted text''; do not use double quotes ("bad form").

• Do not underline. In printing, text is emphasized by changing the type style, usually to slanted or italic type.

• A number of common characters are interpreted as commands, and these must be entered specially, by preceding them with a backslash (\): $ & % # { and } must be typed \$ \& \% \# \{ and \}.

• Authors should refrain from adding vertical or horizontal space. Concentrate on the content of the document and identifying its components with the structural markup commands.

A.2. Math

Mathematical expressions that are part of the running text are delimited by a single dollar sign ($), e.g., $\pi r^2$ yields πr^{2}. To get the appropriately sized superscript or subscript in the roman font, use the \rm command, e.g., $J_{\rm HF}(t)$ produces J_{HF}(t).

Displayed equations can be delimited in several ways. The most concise markup is to bracket the equation between \[ and \] commands, which is equivalent to placing the formula in a displaymath environment. These markup commands will produce unnumbered equations.

Numbered equations can be typeset by typing the formula in an equation environment. A series of related equations that need vertical alignment, e.g., a derivation for which alignment on the equal sign (=) is desired, can be typeset in an eqnarray environment.

While it is possible for authors to assign their own equation numbers, it is easier to let L^{A}TEX number them automatically. By default, L^{A}TEX will number equations sequentially from the beginning of the paper to the end.

It is sometimes appropriate for equation numbering sequences to carry through sections of the paper only. Equations can be numbered (sec-eqn) by placing the command \eqsecnum in the preamble of the document.

A.3. Tables

Tables are notoriously difficult to compose, and great care and patience are usually required before tabular information can be typeset satisfactorily. Tables should be placed in separate table environments, i.e., the tabular material must be enclosed within \begin{table} and \end{table} commands. Tables should have a title or caption and the correct number of descriptive column headings. A single horizontal rule should be set after the column headings with the \tableline command. Do not insert any other horizontal or vertical lines in the body of the table.

The deluxetable environment is intended to make it easier to prepare camera-ready tables, although authors are encouraged to use the deluxetable environment for all tables. Authors who are using the deluxetable styles to prepare camera-ready tables are urged to read the deluxetable section (section 3.13.1) carefully and to look at the many example files in the distribution.

Notes in tables should be marked by \tablenotemark; corresponding text should appear in a \tablenotetext command.

A.4. Cross-referencing

Cross-referencing equations, tables, and figures in text depends upon the use of “keys,” which are defined by the user. The \label command is used to define crossreference keys for L^{A}TEX; \ref is used to refer to them. Keys are simply text strings that serve to label equations, tables, and figures, so that they may be referred to symbolically in the text. Authors should place \label commands immediately after the markup command that starts the structure being referenced. For figure and table captions, place the \label command inside the

\caption, \tablecaption or \figcaption commands,
e.g., \tablecaption{This is a caption.\label{tab1

}}. References to page numbers should not be made.

L^{A}TEX keeps track of autonumbered counters and crossreference information by maintaining an auxiliary file in the same working directory as the source file. The auxiliary file will have an extension of .aux. This file should not be deleted, since subsequent L^{A}TEX processing uses the auxiliary data to resolve references, etc.

The auxiliary file mechanism makes it necessary to run L^{A}TEX on a given source file more than once to ensure that the cross-reference information has been properly resolved. This is most evident when changes are made that affect the number or the placement of equations, tables, and the like. L^{A}TEX will typically issue a warning message that advises the user to “rerun to get cross-references right,” in which case, L^{A}TEX should be run again. If the error message appears after two successive L^{A}TEX runs, it is likely that a reference has been made to an undefined label.

B. AMS Symbols

Authors are permitted to use the American Mathematical Society extra symbol fonts, in addition to the special symbols provided by L^{A}TEX (see below). The amssym substyle file must be included as an argument to the documentstyle command if the AMS symbols are to be used.

     \documentstyle[aasms4,amssym]{article}

The file ams.ps, included with the AASTEX macro package, contains a list of the AMS extra symbol fonts that authors might find useful; this is a PostScript file that can be printed to a local PostScript printer—use ‘lpr’ in Unix—and does not require the AMS symbols to be installed locally.

The AMS symbols are not distributed as part of the AASTEX macro package or as part of the TEX distribution, so authors will need to check with their local TEX administrators to see if these symbols are available for their use. The AMS symbols can be obtained from the American Mathematical Society by anonymous ftp by connecting to ftp.shsu.edu and looking in the di-

rectory tex-archive/fonts/ams. The amssym style file

assumes that the files amssym.def and amssym.tex are in the directory in which the local implementation of TEX looks for input files.

If the AMS fonts and symbols have not been installed locally it is still possible to use them in manuscripts that will be submitted electronically—the local copy of the manuscript will not contain these fonts but the copies produced at the editorial offices and at the Production Offices will be correct.

C. Special symbols

The L^{A}TEX language has a wide variety of special symbols for which markup commands have already been defined. These range from diacritics to exotic mathematical operators.

This section groups L^{A}TEX’s symbols together more or less according to function. Some of these symbols are primarily for use in text; most of them are mathematical symbols and can only be used in L^{A}TEX’s math mode. These tables are excerpted from the L^{A}TEX Command Summary (Botway 1989).

Table C1: Text-mode accents

                                                        ò \`{o}  ō \={o}  oo \t{oo}
                                                        ó \'{o}  ȯ \.{o}  o̧ \c{o}
                                                        ô \^{o}  ŏ \u{o}        o. \d{o}
                                                        ö \"{o}  ǒ \v{o}        ō \b{o}
                                                        õ \~{o}  ő \H{o}

Table C2: National symbols

                                                           œ \oe  å \aa  ł \l
                                                           Œ \OE  Å \AA  Ł \L
                                                           æ \ae  ø \o  ß \ss
                                                           Æ \AE  Ø \O

Table C3: Miscellaneous symbols

                                                       † \dag  § \S        \begin{matrix}c \\ ◯ \end{matrix} \copyright
                                                       ‡ \ddag  ¶ \P  £ \pounds
                                                       # \#  $ \$  % \%
                                                       & \&              \_
                                                       { \{  } \}

Table C4: Math-mode accents

â \hat{a}          ȧ \dot{a}
ǎ \check{a}        ä \ddot{a}
ã \tilde{a}        ă \breve{a}
á \acute{a}        ā \bar{a}
à \grave{a}       a⃗ \vec{a}

Table C5: Greek letters (math mode)

α \alpha  ν \nu
β \beta  ξ \xi
γ \gamma  o o
δ \delta  π \pi
ϵ \epsilon  ρ \rho
ζ \zeta  σ \sigma
η \eta  τ \tau
θ \theta  υ \upsilon
ι \iota  ϕ \phi
κ \kappa  χ \chi
λ \lambda  ψ \psi
μ \mu  ω \omega
ε \varepsilon  ς \varsigma
ϑ \vartheta  φ \varphi
ϱ \varrho
Γ \Gamma  Σ \Sigma
Δ \Delta  Υ \Upsilon
Θ \Theta  Φ \Phi
Λ \Lambda  Ψ \Psi
Ξ \Xi  Ω \Omega
Π \Pi

Table C6: Binary operations (math mode)

± \pm  ∩ \cap
∓ \mp  ∪ \cup
\ \setminus  ⊎ \uplus
· \cdot  ⊓ \sqcap
× \times  ⊔ \sqcup
∗ \ast  ◁ \triangleleft
⋆ \star  ▷ \triangleright
⋄ \diamond  ≀ \wr
∘ \circ  ◯ \bigcirc
• \bullet  △ \bigtriangleup
÷ \div  ▽ \bigtriangledown
< \lhd  > \rhd
∨ \vee  ⊙ \odot
∧ \wedge  † \dagger
⊕ \oplus  ‡ \ddagger
⊖ \ominus  ⨿ \amalg
⊗ \otimes  ≤ \unlhd
⊘ \oslash  ≥ \unrhd

Table C7: Relations (math mode)

≤ \leq  ≥ \geq
≺ \prec  ≻ \succ
⪯ \preceq  ⪰ \succeq
≪ \ll  ≫ \gg
⊂ \subset  ⊃ \supset
⊆ \subseteq  ⊇ \supseteq
⊏ \sqsubset  ⊐ \sqsupset
⊑ \sqsubseteq  ⊒ \sqsupseteq
∈ \in  ∋ \ni
⊢ \vdash  ⊣ \dashv
⌣ \smile  | \mid
⌢ \frown  ∥ \parallel
≠ \neq  ⊥ \perp
≡ \equiv             ≅ \cong
∼ \sim  ▷◁ \bowtie
≃ \simeq  ∝ \propto
≍ \asymp  ⊧ \models
≈ \approx            ≐ \doteq
                       ⋈ \Join

Table C8: Variable-sized symbols (math mode)

∑ ∑ \sum  ⋂ ⋂ \bigcap
∏ ∏ \prod  ⋃ ⋃ \bigcup
∐ ∐ \coprod  ⨆ ⨆ \bigsqcup
∫ ∫ \int  ⋁ ⋁ \bigvee
∮ ∮ \oint  ⋀ ⋀ \bigwedge
⨀ ⨀ \bigodot  ⨂ ⨂ \bigotimes
⨁ ⨁ \bigoplus  ⨄ ⨄ \biguplus

Table C9: Delimiters (math mode)

(  (  )  )
[  [  ]  ]
{  \{  }  \}
⌊  \lfloor  ⌋  \rfloor
⌈  \lceil  ⌉  \rceil
⟨  \langle  ⟩  \rangle
/  /  \ \backslash
|  \vert  ∥  \Vert
↑  \uparrow  ⇑  \Uparrow
↓ \downarrow  ⇓ \Downarrow
↕ \updownarrow  ⇕ \Updownarrow

Table C10: Function names (math mode)

\arccos   \csc    \ker      \min
\arcsin   \deg    \lg       \Pr
\arctan   \det    \lim      \sec
\arg      \dim    \liminf   \sin
\cos      \exp    \limsup   \sinh
\cosh     \gcd    \ln       \sup
\cot      \hom    \log      \tan
\coth     \inf    \max      \tanh

Table C11: Arrows (math mode)

                                                 ← \leftarrow  ⟵ \longleftarrow
                                                 ⇐ \Leftarrow  ⟸ \Longleftarrow
                                                 → \rightarrow  ⟶ \longrightarrow
                                                 ⇒ \Rightarrow  ⟹ \Longrightarrow
                                                 ↔ \leftrightarrow  ⟷ \longleftrightarrow
                                                 ⇔ \Leftrightarrow  ⟺ \Longleftrightarrow
                                                 ↦ \mapsto  ⟼ \longmapsto
                                                 ↩ \hookleftarrow  ↪ \hookrightarrow
                                                 ↼ \leftharpoonup  ⇀ \rightharpoonup
                                                 ↽ \leftharpoondown  ⇁ \rightharpoondown
                                                 ⇀ ↽ \rightleftharpoons ⇝ \leadsto
                                                 ↑ \uparrow  ⇕ \Updownarrow
                                                 ⇑ \Uparrow  ↗ \nearrow
                                                 ↓ \downarrow  ↘ \searrow
                                                 ⇓ \Downarrow  ↙ \swarrow
                                                 ↕ \updownarrow  ↖ \nwarrow

Table C12: Miscellaneous symbols (math mode)

                                                      ℵ \aleph  ′ \prime
                                                      h̄ \hbar  ∅ \emptyset
                                                       ı \imath  ∇ \nabla
                                                       ȷ \jmath  √ \surd
                                                       ℓ \ell  ⊤ \top
                                                      ℘ \wp  ⊥ \bot
                                                      ℜ \Re  ∥ \|
                                                      ℑ \Im           ^{}\angle
                                                      ∂ \partial  △ \triangle
                                                      ∞ \infty  \ \backslash
                                                      □ \Box  ◇ \Diamond
                                                      ∀ \forall  ♯ \sharp
                                                      ∃ \exists  ♣ \clubsuit
                                                      ¬ \neg  ♢ \diamondsuit
                                                       ♭ \flat  ♡ \heartsuit
                                                       ♮ \natural  ♠ \spadesuit
                                                      ℧ \mho